### v2

# FiRST: Finetuning Router-Selective Transformers for Input-Adaptive Latency Reduction

Abstract
**footnotetext: Equal Contribution22footnotetext: Work done during internship at Adobe Research
Auto-regressive Large Language Models (LLMs) demonstrate remarkable performance across different domains such as vision and language processing. However, due to sequential processing through a stack of transformer layers, autoregressive decoding faces significant computation/latency challenges, particularly in resource-constrained environments like mobile and edge devices. Existing approaches in literature that aim to improve latency via skipping layers have two distinct flavors - 1) Early exit, and 2) Input-agnostic heuristics where tokens exit at pre-determined layers irrespective of input sequence. Both the above strategies have limitations - the former cannot be applied to handle KV Caching necessary for speed-ups in modern framework and the latter does not capture the variation in layer importance across tasks or more generally, across input sequences. To address both limitations, we propose FiRST, an algorithm that reduces inference latency by using layer-specific routers to select a subset of transformer layers adaptively for each input sequence - the prompt (during the prefill stage) decides which layers will be skipped during decoding. FiRST preserves compatibility with KV caching enabling faster inference while being quality-aware. FiRST is model-agnostic and can be easily enabled on any pre-trained LLM. Our approach reveals that input adaptivity is critical - indeed, different task-specific middle layers play a crucial role in evolving hidden representations depending on tasks. Extensive experiments show that FiRST significantly reduces latency while outperforming other layer selection strategies in quality metics. It retains competitive performance to base model (without layer skipping) and in some cases, even improves upon it. FiRST is thus a promising and efficient solution for LLM deployment in low-resource environments.
1 Introduction
Large Language Models (LLM’s) have revolutionized the fields of Natural Language Processing and Computer Vision achieving incredible performance on a diverse set of benchmark tasks. However, the scale of these LLM’s characterized by billions of parameters hinder their adoption in resource-constrained environments with memory, latency and compute being the main challenges. In this work, we focus on the latency aspect which becomes the most significant bottleneck for tasks such as machine translation, question answering, summarization particularly on devices, such as laptops and mobile phones. As noted by (Schuster et al., 2022), the auto-regressive nature of decoding in LLM’s further pronounces the latency bottleneck.
Transformer based LLMs have several stacks of layers (including attention and FFN layers) leading to high latency and compute requirements, making inference very slow or even infeasible in resource constrained settings. This is because of the sequential processing of tokens through all the layers for every input sequence and task. However, it is important to note that in the real world, there is a lot of heterogeneity in input sequences and tasks. (Schuster et al., 2022; Sun et al., 2022) noted that the generations made by LLMs can have varying levels of difficulty and certain generations can be solved with reduced compute, by exiting the transformer stack early. At the same time, it has been noted in recent works (Wendler et al., 2024) that inference forward pass proceeds in phases through the layers of transformer based models, with different types of information being extracted or mapped at different phases (sequences of layers) for certain tasks such as translation. Motivated by these and other related works, we hypothesize that different sequential combinations of layers are important for different input sequences and tasks. Learning the right sequential combination of layers can help reduce inference latency and compute for on-device scenarios. However, there are several challenges. Any algorithm for determining the “right” combination of layers should minimize any quality loss, be compatible with other latency reduction strategies such as KV cache handling and be learnable with minimal compute/training overhead.
In the last few years, several promising approaches have been proposed in literature that adaptively prune layers at each decoding step. Token-level early exit proposed in (Schuster et al., 2022; Sun et al., 2022) allow tokens to exit the transformer layer stack early based on different strategies to compute the confidence or saturation level. (Elhoushi et al., 2024; Elbayad et al., 2020; Zhang et al., 2019) extended this idea to incorporate layer skipping at a token level during training. While token level early exit is a useful idea in theory, it suffers from a major limitation of incompatible KV caching in practice (Del Corro et al., 2023). The incompatibility stems from having to recompute KV caches for preceding tokens if we have a delayed exit point for latter tokens, often resulting in loss of early exit advantages. This limits its practical adoption since KV cache is crucial in significantly speeding up auto-regressive decoding.
Recently, (Liu et al., 2024; Del Corro et al., 2023; Song et al., 2024) have proposed input-agnostic layer skipping at token level, that handle KV cache appropriately as well as retain the advantage of adaptive partial computation. In these solutions, tokens exit at pre-determined layers irrespective of the input sequence, and for all sequences in a batch, tokens at the same position in a sequence exit at the same layer. Furthermore, tokens at latter parts of the sequence are constrained to exit earlier than the previous tokens to ensure that there is no redundant KV cache re-computation. These solutions are heuristic based and impose hard rules and constraints irrespective of input sequences, which can lead to drop in output quality. Others (Jaiswal et al., 2024; Chen et al., 2024) have proposed skipping layers by identifying redundant ones through computing cosine similarity of (input/output) representations of a layer. However, their strategy does not take into account that several middle layers are crucial (see (Liu et al., 2024)) and furthermore, final prediction capability of full model is not taken into account while deciding which layers to skip. Importantly, in none of the works described above, the strategy of selecting layers for skipping is sequence dependent. Furthermore, they do not consider finetuning the models in a way such that not only the performance improves but also the model learns to skip layers appropriately.
Our goal is to design an input-adaptive learnable layer selection strategy with quality aware latency gains that is also able to handle the KV cache appropriately. Ideally, for every input sequence and task, we want to predict the optimal (sequential) combination of layers at inference time, such that quality loss is minimum and the latency gains are as high as possible. We want to do this with expending very little compute/additional training and appropriate handling of KV cache. We propose to do this via training routers. Based on the output of each layer for a sequence, a router will decide whether or not to skip the subsequent layer in the transformer architecture. Since the decision is at a sequence level, KV cache issues do not arise, as all tokens in a sequence would pass through the same set of layers. Finally, we fine-tune the model combined with trained routers using LoRA adapters to improve the quality significantly while retaining the latency gains. As an added bonus, LoRA finetuning smoothens the layer skipping and further highlights the varied importance of layers based on input sequence. We summarize our contributions:
  1. 1.
    We propose a training and inference algorithm FiRST that incorporates layer-specific routers for selecting layers in an input-adaptive manner. The layer selection is uniform for all tokens in a sequence, thus handling KV caches without introducing additional compute/latency. FiRST can be applied on top of any pre-trained model.
  2. 2.
    We propose LoRA based finetuning on top of router based layer selection to improve quality while retaining latency gains. This also smoothens layer selection.
  3. 3.
    Finally, we demonstrate an extensive set of experiments with FiRST on multiple datasets for 3 distinct tasks namely Machine Translation, Summarization and Question Answering, and 2 different open model architectures namely LLaMA-3-8B and LLaMA-3.2-3B. We show that for the same target speed-up, FiRST significantly improves performance across tasks as compared to baselines and prior works.
Due to space constraints, we delegate a study of other related works and orthogonal approaches (for e.g. model compression) for exploring latency/performance tradeoff to Appendix A.1.
2 Problem Statement
Our goal is to exploit the heterogeneity in inputs and tasks to selectively use LLM layers in a quality-aware manner for reducing inference latency and compute for on-device constraints. Ideally, we want to select an optimal sub-sequence of layers within a transformer architecture for a given input and task, such that the overall latency, as well as expended computation, are both low, while quality is comparable to the un-modified case where every input sequence passes through every layer. For ease of explanation, without loss of generality, we assume the task is same and simply consider an input sequence for describing the problem.
Let us consider an an input sequence $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$ with $n$ tokens. Let there be $m$ transformer layers in the model, where the $i^{th}$ transformer layer is represented as the function $\phi_{i}()$. As stated lucidly in (Wendler et al., 2024), $\mathcal{X}$ is first converted to an initial latent representation $\mathcal{H}_{0}=\{H^{1}_{0},H^{2}_{0},\ldots,H^{n}_{0}\}$, where $H^{0}_{j}\in\mathbb{R}^{D},\forall j\in[n]$ is a look-up from a learned embedding dictionary corresponding to the $j^{th}$ token. Thereafter, every transformer layer $\phi_{i}()$ operates on the latent vectors $\mathcal{H}_{i}$ to generate the embedding for the $i^{th}$ layer as follows. For the $j^{th}$ token,
$H^{j}_{i}$ $=H^{j}_{i-1}+\phi_{i}(H^{1}_{i-1},H^{2}_{i-1},\ldots,H^{j}_{i-1})$ (1)
Let the (gold) output or generated sequence for an input sequence $\mathcal{X}$ that passed through all $m$ layers of the model with full computation be $\mathcal{Y}^{*}_{\mathcal{X}}$. Our hypothesis is that for a given input sequence (and task), there exists an optimal subsequence of functions $\mathcal{F}_{OPT}(\mathcal{X})$ out of the full sequence $\{\phi_{i},i\in[m]\}$ such that the output generated by passing through this subsequence: $\mathcal{Y}_{OPT,\mathcal{X}}\approx\mathcal{Y}^{*}_{\mathcal{X}}$. More formally, if $Q$ is a quantitative quality measure on $\mathcal{Y}$, and $\epsilon\rightarrow 0$ is tolerance in deviation in quality from the gold output, then we hypothesize that there exists an optimal subsequence, using the minimum number of layers, $\mathcal{F}_{OPT}(\mathcal{X})$, such that:
$Q\left(\mathcal{Y}_{OPT,\mathcal{X}}\right)$ $\geq(1-\epsilon)Q\left(\mathcal{Y}^{*}_{\mathcal{X}}\right),\forall\mathcal{X}.$ (2)
The optimality above is with respect to the minimum subsequence of layers that can help achieve the above, to minimize latency while keeping quality unaffected. Note that, the optimal subsequence $\mathcal{F}_{OPT}(\mathcal{X})$ need to be obey the same autoregressive computation on previous tokens as given in Equation 1. Hence, any algorithm that determines the optimal subsequence, need to be compatible with KV cache handling, to avoid the re-computation of values for tokens preceding the current token.
The potential number of subsequences for $m$ layers is $2^{m}$, hence a brute force is infeasible and also beats the purpose of such a layer selection in the first place: reducing latency and compute. In the absence of any known substructure in the behaviour of the latent layers on each input sequence, it is difficult to arrive at the optimal solution polynomially or with low additional latency or compute.
We propose to learn an approximation of the optimal subsequence of layers for any input sequence with low additional latency and minimal training.
3 Proposed Solution: FiRST
Let us first understand what it entails to learn an optimal subsequence of layers for any input. Consider the full transformer sequence to be =$\mathcal{F}^{*}=\{\phi_{1},\phi_{2},\ldots,\phi_{m}\}$. Any optimal subsequence for an input $\mathcal{X}$: $\mathcal{F}_{OPT,\mathcal{X}}$ could be thought of as finding an optimal path through a binary tree of functions. Formally, let every level in the binary tree correspond to a transformer layer and the $0^{th}$ layer corresponds to the initial embedding look up; i.e., at depth $i\in[m]$, there would be $2^{i}$ nodes, each corresponding to either $\phi_{i}$ or $\overline{\phi_{i}}$, where the former denotes that a particular transformer layer is included in the optimal path whereas the latter denotes that it is not included. Each (of the $2^{i-1}$ nodes) $\phi_{i}$ or $\overline{\phi_{i}}$ has two children, corresponding to the next transformer layer: $\phi_{i+1}$ and $\overline{\phi_{i+1}}$ (See Figure 1). In such a tree structure, for example, the path $\{\phi_{i},\overline{\phi_{i+1}},\phi_{i+2}\}$ indicates the subsequence of transformer layers $\{\phi_{i},\phi_{i+2}\}$. For any transformer layer $\phi_{i}$ in this tree, let $Anc(\phi_{i})=k,0\leq k<i$ denote the the lowest ancestor node where the corresponding transformer node $\phi_{k}$ is included in the sequence. In the above example, $Anc(\phi_{i+2})=\phi_{i}$.
Figure 1: Binary Tree representation of layer selection.
Consider a sequence of functions $\mathcal{F}$, where for level $i$, $Anc(\phi_{i})=\phi_{k}$. The autoregressive computations for the $j^{th}$ token in the input sequence (originally Eq 1), would now be modified as:
$H^{j}_{i}$ =$=\begin{cases}H^{j}_{k},&\text{if }\phi_{i}\notin\mathcal{F},\\ H^{j}_{k}+\phi_{i}(H^{1}_{k},H^{2}_{k},\ldots,H^{j}_{k}),&\text{if }\phi_{i}\in\mathcal{F}.\end{cases}$ (3)
Our problem translates to navigating this binary tree to find the optimal path $\mathcal{F}_{OPT}$ for an input sequence and task. Since there are $2^{m}$ paths in this tree, we propose to approximate the optimal by making a decision in a greedy fashion at each node. Formally, we add a (lightweight and fast) router $R_{i}$ before every transformer layer $\phi_{i}$ in the model, that will predict whether $\phi_{i}$ will be selected or not.
Our aim is to learn to predict the layer choice at a sequence level (not token) to maintain compatibility with the autoregressive computations and avoid re-computation of of KV cache values. Moreover, we should spend minimal compute for learning the $R_{i}$ functions. Finally, $R_{i}$ functions should not add any significant latency to the overall computation.
FiRST modifies any off-shelf pre-trained transformer based model by incorporating and training a router or probability function $R_{i}$ before every transformer layer $\phi_{i}$. The output of $R_{i}$ is a score $\rho_{i}$ denoting the probability of selecting $\phi_{i}$ in the layer sequence. During inference, $\rho_{i}$ is rounded to determine selection of $\phi_{i}$. Let $\lfloor\rho_{i}\rceil=1\text{ if }\rho_{i}\geq 0.5,\text{ else }0$. Equation 1 is now modified as:
$H^{j}_{i}$ $=H^{j}_{i-1}+\lfloor\rho_{i}\rceil\cdot\phi_{i}(H^{1}_{i-1},H^{2}_{i-1},\ldots,H^{j}_{i-1})$
This recursively approximates Eq 3 for the optimal $\mathcal{F}$ in a probabilistic, greedy manner. We train the functions $R_{i}$ on datasets and tasks, and further fine tune using LoRA adapters to make the layer selections smooth and improve the output quality.
4 FiRST Framework and Algorithm
In this section, we describe the training and inference frameworks for FiRST in details. We discuss how to train Routers to be adaptive to input sequences. Given an off-the-shelf pre-trained LLM, we propose two training phases. In the first phase, we train a router for each layer that decides whether the input sequence should skip the layer. In the second phase, to tackle the issue of unseen skipping during pre-training, we fine-tune the router-augmented LLM keeping router weights fixed to ensure the model improves performance on the target dataset without reducing the skipping level.
4.1 Adaptive Router Module
The adaptive router module is a single-layer neural network without bias, positioned before every layer in the model. During training of the router, all model parameters except the router weights remain frozen. For the first layer, it takes the tokenized input, and for each of the subsequent layers, it takes the output of the preceding layer as input. Mathematically speaking, for any layer $i$, given a batch of $B$ tokenized inputs sequences, where each sequence has $n$ tokens and is embedded in to $\mathbb{R}^{D}$, the adaptive router module takes as input a $B\times n\times D$ tensor output of layer $(i-1)$ and outputs a $B\times n\times 1$ tensor. Subsequently, corresponding to each value (or, token) in the $B\times n\times 1$ tensor, we apply a sigmoid function to ensure that all entries in the tensor are in the interval $[0,1]$. Following this, we take a mean operation at the sequence level - we take a mean of all the weights in a sequence to output a $B\times 1\times 1$ tensor. For each sequence in the batch, the corresponding entry is the probability $\rho_{i}$ with which the sequence passes through the layer $i$. The input sequence skips the layer $i$ with probability $1-\rho_{i}$. During training, the output of a layer is modified using a skip connection, incorporating the probability $\rho_{i}$ (see Figure: 2).
The routers are trained to encourage skipping by reducing the probabilities $\{\rho_{i}\}_{i}$ using a regularizer, to approximate the optimal subsequence for minimizing the latency. The training task is modeled as a language modeling task, specifically next token prediction. The loss function comprises of 3 terms:
  • Cross-entropy loss: Standard difference between actual and predicted probability distributions to ensure the quality of generation: $\mathcal{L}_{\mathrm{CE}}=-\sum_{x\in\mathcal{X}}\mathcal{Y}^{*}_{\mathcal{X}}\log(\hat{\mathcal{Y}})$.
  • Regularization loss: Adds a penalty term to reduce overfitting to noise: $\mathcal{L}_{\mathrm{Reg}}=\sum_{i\in[m]}||R_{i}||^{2}$, where $||R_{i}||^{2}$ denotes the $\ell_{2}$ norm of the router weights for the $i^{\text{th}}$ layer router, and there are $m$ layers in the model.
  • Non-skip penalization loss: This is the summation of probability values across all layers of the model architecture: $\mathcal{L}_{\mathrm{PP}}=\sum_{i\in[m]}\rho_{i}$
The total loss $\mathcal{L}$ is a linear combination of these three terms: $\mathcal{L}=\mathcal{L}_{\mathrm{CE}}+\lambda\cdot\mathcal{L}_{\mathrm{Reg}}+\alpha\cdot\mathcal{L}_{\mathrm{PP}}$, where $\alpha$ manages the tradeoff between quality and latency.
4.2 LoRA Compensation Module
Skipping layers naturally leads to some performance loss - especially so since the pre-trained model was not trained to skip layers. To compensate for the loss in performance caused by skipping layers, we finetune the router-augmented pre-trained model on the downstream task 111similar to Quantization Aware Training such as QLoRA Dettmers et al. (2024) - compensates for model compression using Low Rank Adapters (LoRA). During finetuning, the router parameters are frozen while trainable LoRA adapters are added to both the FFN (Feed-Forward Network) and the attention modules of each layer of the pre-trained model. In order to maintain the skipping level, we again add a non-skip penalization loss component during finetuning with scaling hyper-parameter $\beta$. This is essential even though the router weights are frozen because standard finetuning alters the hidden representations of the input sequence in a manner such that no layers are skipped. Note that the LoRA adapters do not lead to any latency overhead during inference.
Skip (%) Model Type English-to-German English-to-Chinese
BLEU-1 BLEU-2 COMET BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 41.78 21.74 93 56.94 35.56 82.66
Base 37.17 18.57 87.13 38.02 22.46 68.95
15 Skip Decode Router + LoRA 23.04 10.52 55.62 28.73 15.84 55.98
Router 3.99 1.18 23.33 4.74 2.33 21.75
Random Skip Router + LoRA 30.43 10.98 66.25 47.88 25.75 67.32
Router 26.54 8.77 60.27 36.60 18.66 59.89
Unified Skip Router + LoRA 28.92 10.64 59.34 46.61 25.01 69.58
Router 23.23 7.85 59.26 27.28 13.35 54.57
FiRST (Ours) Router + LoRA 38.01 17.89 82.14 48.35 26.57 68.63
Router 28.83 11.8 67.74 17.55 8.68 42.76
25 Skip Decode Router + LoRA 13.67 6.00 31.47 20.03 10.85 33.85
Router 3.24 0.92 21.55 3.78 1.84 20.93
Random Skip Router + LoRA 6.01 0.91 29.71 11.69 4.66 27.73
Router 3.65 0.49 29.95 7.37 2.81 35.16
Unified Skip Router + LoRA 15.67 3.36 31.69 34.90 15.75 50.59
Router 12.58 2.65 32.15 17.74 7.35 38.74
FiRST (Ours) Router + LoRA 17.84 4.14 34.95 35.79 15.66 56.92
Router 9.67 1.37 26.01 11.01 3.23 25.45
Table 1: Machine Translation Results for LLaMA-3-8B: BLEU-1, BLEU-2, and COMET scores are reported for English-to-German and English-to-Chinese tasks across varying skip levels. FiRST consistently achieves the highest BLEU-1 and BLEU-2 at 15% skipping, for both translation directions. Similarly, high BLEU-1 and COMET scores are obtained for 25% skipping rate.
4.3 Inference for FiRST
During inference, for the input sequence, each router (corresponding to a layer) outputs a number in the interval $[0,1]$. If this number is greater than or equal to 0.5, the sequence passes through the layer. Otherwise, the sequence skips the layer (Fig. 2).
Figure 2: Skip connection used for router training. With probability $p$, the sequence is processed by the layer and with probability $1-p$, the layer is skipped. During inference, routers make the decision of whether a sequence will skip a particular layer or pass through it.
Below, we discuss some salient points about the functioning of the router during inference to handle KV Cache appropriately:
  1. 1.
    Prefill phase handling: Skipping is not allowed during prefill phase. This ensures the first token is generated correctly, which is crucial for WMT tasks, as they are highly sensitive to the correct generation of the first token in the target language. It has been observed in prior works (Liu et al., 2024) that skipping during prefill phase is detrimental to performance during inference.
  2. 2.
    Fixed router decisions during decoding and handling KV Cache: During the prefill phase, the decisions made by the routers are cached. During the decoding phase, every token adheres to the cached decision made during prefill. In other words, for a particular layer, if a router outputs a number less than 0.5 during prefill, the number is fixed for the decoding steps and therefore the same layer will be skipped by all tokens during decoding. Similarly, if the router outputs a number more than 0.5 during prefill, the same layer will be processing all tokens during decoding. Such a step ensures that for each decoding step and each layer that is not skipped, the KV cache for all previous tokens is available for that layer. This approach effectively addresses the caching issues encountered in early exit strategies, ensuring consistent decisions across the decoding process.
5 Experiments
Skip (%) Model Type English-to-German English-to-Chinese
BLEU-1 BLEU-2 COMET BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 37.57 17.43 89.72 51.81 30.04 79.13
Base 31.19 13.67 81.66 32.10 17.92 61.84
15 Skip Decode Router + LoRA 23.24 11.07 44.58 38.14 21.68 46.70
Router 9.93 3.89 32.74 9.59 4.84 34.14
Random Skip Router + LoRA 18.99 4.78 47.26 38.75 17.39 57.79
Router 12.11 2.77 36.30 13.56 5.64 35.86
Unified Skip Router + LoRA 21.65 5.93 44.72 36.96 17.35 57.10
Router 16.41 3.99 39.81 22.43 9.37 45.16
FiRST (Ours) Router + LoRA 24.68 7.13 60.29 45.66 23.66 67.45
Router 16.47 4.11 43.04 22.69 10.92 54.55
25 Skip Decode Router + LoRA 16.85 7.62 32.33 27.81 15.74 42.01
Router 9.14 3.51 27.64 7.01 3.41 29.51
Random Skip Router + LoRA 8.95 1.07 30.97 25.12 9.85 45.53
Router 5.32 0.69 27.22 10.07 3.69 31.50
Unified Skip Router + LoRA 15.14 2.65 39.81 30.56 12.27 42.30
Router 10.21 1.62 30.86 15.21 5.51 31.63
FiRST (Ours) Router + LoRA 18.89 4.06 45.38 32.92 13.74 41.66
Router 9.51 1.44 29.08 10.46 3.55 27.83
Table 2: Machine Translation Results for LLaMA-3.2-3B: BLEU-1, BLEU-2, and COMET scores are shown for English-to-German and English-to-Chinese tasks at different skip levels. FiRST consistently achieves the highest BLEU-1 and COMET scores for English-to-German at both skipping percentages and highest BLEU-1 for English-to-Chinese.
Skip (%) Model Type EM F1
0 Original Model Base + LoRA 73.93 85.99
Base 19.46 36.73
10 Skip Decode R + LoRA 60.14 65.33
Router 16.38 31.48
Random Skip R + LoRA 65.73 80.08
Router 18.25 33.75
Unified Skip R + LoRA 55.54 74.58
Router 17.39 32.91
FiRST (Ours) R + LoRA 70.85 83.61
Router 14.58 31.52
20 Skip Decode R + LoRA 45.00 55.10
Router 10.68 26.69
Random Skip R + LoRA 47.79 66.37
Router 6.71 22.46
Unified Skip R + LoRA 52.87 69.28
Router 18.18 32.51
FiRST (Ours) R + LoRA 60.60 75.49
Router 13.21 27.48
Skip (%) Model Type EM F1
0 Original Model wLoRA 73.07 84.17
Base 18.92 37.74
10 Skip Decode R + LoRA 60.79 75.00
Router 20.00 31.55
Random Skip R + LoRA 64.78 77.27
Router 13.76 28.59
Unified Skip R + LoRA 65.03 77.53
Router 13.16 32.31
FiRST R + LoRA 69.44 81.35
Router 12.79 28.37
20 Skip Decode R + LoRA 40.12 40.00
Router 20.45 37.62
Random Skip R + LoRA 11.32 38.34
Router 6.75 15.51
Unified Skip R + LoRA 37.39 52.49
Router 7.81 18.20
FiRST R + LoRA 39.70 54.59
Router 5.52 15.33
Table 3: Quality Analysis on Question Answering (SQuAD v1.1): LLaMA-3-8B (left) and LLaMA-3.2-3B (right). Exact Match (EM) and F1 scores are reported for varying skipping levels. Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning.
Skip (%) Model Type BERT R-1 R-L
0 Original Model wLoRA 84.87 28.46 16.99
Base 82.29 23.49 14.66
15 Skip Decode R + LoRA 84.74 22.04 17.54
Router 82.53 13.68 9.30
Random Skip R + LoRA 83.70 24.60 15.01
Router 81.10 19.64 13.07
Unified Skip R + LoRA 84.25 24.35 14.3
Router 80.3 16.61 10.95
FiRST (Ours) R + LoRA 85.14 31.8 20.13
Router 81.25 20.2 13.01
27 Skip Decode R + LoRA 79.92 10.67 10.32
Router 77.27 9.59 7.00
Random Skip R + LoRA 76.40 11.45 7.89
Router 77.45 12.56 9.08
Unified Skip R + LoRA 80.28 15.94 9.89
Router 77.43 10.97 7.68
FiRST (Ours) R + LoRA 77.5 14.65 10.45
Router 75.6 9.39 6.92
Skip (%) Model Type BERT R-1 R-L
0 Original Model wLoRA 84.89 28.37 17.02
Base 71.85 19.34 12.00
15 Skip Decode R + LoRA 83.20 21.71 13.74
Router 80.97 9.74 6.87
Random Skip R + LoRA 79.52 20.18 12.10
Router 68.20 10.10 7.10
Unified Skip R + LoRA 81.53 18.89 11.72
Router 70.01 12.49 8.68
FiRST (Ours) R + LoRA 83.17 26.47 16.79
Router 70.98 16.47 10.51
24 Skip Decode R + LoRA 78.55 15.83 6.74
Router 76.91 13.29 8.86
Random Skip R + LoRA 80.00 16.33 10.07
Router 67.88 8.49 5.96
Unified Skip R + LoRA 79.31 15.88 10.69
Router 68.86 9.17 6.97
FiRST (Ours) R + LoRA 80.25 21.28 13.89
Router 69.17 12.36 8.43
Table 4: Quality Analysis on Summarization (CNN/DM dataset) on LLaMA-3-8B (left) and LLaMA-3.2-3B (right): BERT F1, Rouge-1 and Rouge-L scores are reported for varying skipping levels. Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning.
We conduct experiments on three benchmark tasks namely Machine Translation, Text Summarization and Question Answering demonstrating both robustness and scalability of FiRST.
Datasets: For machine translation, we use WMT development sets (2017–2020) for English-to-Chinese and English-to-German tasks, evaluating performance on the WMT 2022 test set, which covers diverse domains such as news, social media, e-commerce, and conversational contexts. For summarization, we use the CNN/DailyMail dataset, with 4000 randomly selected training samples and evaluation on the standard test set of 11,490 samples. In Question Answering, we utilize SQuAD v1.0, training on 4000 randomly selected samples from the 100k question-answer pairs and evaluating on the validation set, as test set labels are unavailable. Appendix A.2 contains detailed descriptions.
Evaluation Metrics: We use standard metrics to evaluate the quality of generated output in each of the three tasks. For Machine Translation, we benchmark using BLEU scores and COMET where the latter provides a more nuanced assessment beyond $n$-gram used in BLEU. For Summarization, we use ROUGE scores and BERT Score - as before, the latter captures meaning-based similarity. For Question Answering, we use Exact Match and F1 Score as the metrics to benchmark the output quality. Finally, for benchmarking latency, we look at the TPOT (Time Per Output Token). TPOT evaluates the average time taken to produce each output token and is calculated for GPU to gauge overall decoding performance. Appendix A.3 contains a detailed description of evaluation metrics. Additionally, hyper-parameters used during training and inference can be found in Appendix A.4.
5.1 Baselines for comparison
We report the latency improvement and quality numbers relative to the base models (no skipping).
  • Random Skipping: We skip a set of $k$ layers randomly where $k$ depends on the target speedup.
  • Skip Decode: We implement Skip Decode Del Corro et al. (2023) method that features a monotonic decrease in processing layers, enabling later tokens to leverage the computational resources used for earlier ones.
  • Unified Skipping: This, to the best of our knowledge, is the state-of-the-art method that relies on using a heuristic-based strategy for retaining layers at fixed intervals. We replicate the algorithm in (Liu et al., 2024) and compare performance both with and without LoRA fine-tuning across various skipping percentages.
5.2 Experimental Results on Different Tasks
WMT: Our experiments demonstrate that FiRST performs consistently well in different machine translation tasks (EN-to-DE, EN-to-ZH) across different models. For LLaMA-3.8B, for $15\%$ skipping, FiRST achieves a latency improvement of 10-12% on TPOT (see Tables 1 and 5) while being competitive to the base model + LoRA (gold reference) in quality. In most cases, it significantly outperforms other layer skipping strategies (Random, Skip Decode, Unified Skipping) and in other cases, it is comparable in quality. In comparison to the gold output (Base + LoRA), FiRST is $\geq 80\%$ in COMET scores (semantic metric), $\geq 85\%$ in BLEU-1 scores, and $\geq 75\%$ in BLEU-2 scores (syntax metrics) for both EN-to-DE and EN-to-ZH translations. For $25\%$ skipping, FiRST achieves significant improvement in quality over other strategies, in almost all metrics, while achieving $\sim 18\%$ reduction in TPOT.
For LLaMA-3.2-3B, the latency improvement is $\sim 10\%$, with quality scores being significantly higher than other layer skipping strategies (Table 6). It is within $65-85\%$ of BLEU-1 and COMET scores (Table 2) for EN-DE and EN-ZH of the gold (LoRA fine-tuned base model).
CNN/DailyMail Dataset: For LLaMA-3-8B, at roughly 15% skipping level, our method outperforms the base model + LoRA (Table: 4) while obtaining a 12% improvement in TPOT (Table 5). For LLaMA-3.2-3B, at $15\%$, the quality is comparable ($\sim 98\%$) to gold and other baselines with $12\%$ improvement in TPOT. At $24\%$, FiRST is significantly better than other layer skipping strategies, while achieving $>20\%$ improvement in latency.
Model Type $\sim$ Skipping (%) English-to-German English-to-Chinese
TPOT TPOT
Base + LoRA 0 1x 1x
R + LoRA 15 0.90x 0.88x
R + LoRA 25 0.82x 0.83x
R + LoRA 35 0.69x 0.68x
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.88x
R + LoRA 20 0.81x
R + LoRA 27 0.76x
Table 5: TPOT variation of LLaMA-3-8B on WMT (left) and CNN/DM (right) for FiRST. These values are relative to the LoRA fine-tuned base model. Fine-tuning improves TPOT and quality significantly.
Model Type $\sim$ Skipping (%) English-to-German English-to-Chinese
TPOT TPOT
Base + LoRA 0 1x 1x
R + LoRA 15 0.90x 0.91x
R + LoRA 25 0.78x 0.75x
R + LoRA 35 0.69x 0.74x
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.88x
R + LoRA 24 0.79x
R + LoRA 28 0.77x
Table 6: TPOT variation of LLaMA-3.2-3B on WMT (left) and CNN/DM (right) for FiRST. These values are relative to the LoRA fine-tuned base model. Fine-tuning improves TPOT and quality significantly.
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.94x
R + LoRA 24 0.83x
R + LoRA 28 0.74x
[…]
TPOT
Base + LoRA 0 1x
R + LoRA 15 0.94x
R + LoRA 24 0.84x
R + LoRA 28 0.72x
Table 7: TPOT variation of LLaMA-3-8B (left) and LLaMA-3.2-3B (right) on SQuAD dataset for FiRST. These values are relative to the LoRA fine-tuned base model. Fine-tuning improves TPOT and quality significantly.
SQuAD Dataset: For the LLaMA-3-8B model, FiRST is $>95\%$ in output quality of gold (base + LoRA) (Table 3), with overall latency gains of 6-16% (Table 7). It is significantly better in quality than all other baselines across all metrics for different levels of skipping. For LLaMA-3.2-3B, again FiRST is $>95\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 3) with gains in latency of 6-16% overall (Table 7) over the LoRA fine-tuned base model. Moreover, it is better than all other layer skipping strategies across all metrics. Detailed results for an additional skipping percentage are provided in Appendix A.8.
Layer-wise Skipping Patterns: Layer-wise skipping varies significantly across tasks, reflecting the task-specific importance of each layer. For LLaMA-3-8B at a 15% skipping rate, layers 7–9 and 21 are fully skipped in English-to-German, with partial skipping in layer 18. In English-to-Chinese, layers 7–9, 16, and 21 are fully skipped, while layer 20 is partially skipped. For Summarization, layers 20, 22, and 23 are fully skipped, with partial skipping in layers 19, 21, and 26. Some layers are skipped less than 10%, indicating their necessity for specific sequences. This task-specificity is also evident in Question-Answering, where only layer 22 is fully skipped, and skipping patterns depend on the input. Detailed statistics are in Appendix A.7.
Computational overhead of Routers: We also examine the computational overhead introduced by training the routers. This overhead remains minimal, as the router parameters are significantly smaller than the total number of trainable parameters in each model (0.0027% for LLaMA-3-8B and 0.0016% for LLaMA-3.2-3B). The proportion of the total computational time spent on the router operations compared to the entire forward pass of the model is close to 0.3% for both models suggesting their overhead does not significantly affect the efficiency of the model during inference or training.
6 Conclusion
We propose a new framework FiRST for layer selection corresponding to input sequence and task towards reducing latency in a quality aware manner. This is sequence dependent and operates in a KV cache compatible manner. For an optimal skipping rate of around 15%, FiRST achieves 10-20% reduction in latency while being quality neutral (approximately 80% or more in quality metrics compared to base model) on multiple tasks/model architectures such as Machine Translation, Summarization and Question Answering, on well known open source datasets, and some cases, even improving upon base model, while significantly outperfoming other layer selection strategies on most quality metrics.
7 Limitations
FiRST algorithm selects layers in a greedy, myopic way one layer at a time, corresponding to sequences (and tasks). A more optimal way of doing this would be to estimate a subsequence of layers to traverse through instead of one layer at a time. We intend to address this in future work. We would like to select a more optimal subset (subsequence) of layers which will increase the output quality while reducing latency even further.
8 Ethical Concerns
There are no ethical concerns to the best of our knowledge.
Appendix A Appendix
A.1 Related Work
Early Exit:
Several works have been proposed in the early exit theme (Zhu, 2021; Zhou et al., 2020; Xin et al., 2020; Liu et al., 2020; Li et al., 2020; Hou et al., 2020; Schuster et al., 2022) where adaptive compute is used for different parts of the token sequence. While these approaches have been popular for encoder-only models which processes the entire sequence as a whole, they have faced challenges in generation tasks. The main limitation of these set of techniques are their inability to handle KV caching appropriately which is crucial for multi-fold speed-ups in current LLM architectures. We emphasize that in our work, we assign varying compute to sequences in different batches but within the same sequence, we assign the same compute to every token.
Input Agnostic Heuristics:
In Skip Decoding (Del Corro et al., 2023), initial tokens pass through more layers than later ones, contradicting the observation that later tokens are harder to decode (Liu et al., 2024). Additionally, Skip Decoding skips several bottom layers for most tokens, causing undesirable sub-network imbalance. To address this, Unified Layer Skipping (Liu et al., 2024) proposes a discrete skipping strategy that is uniform for all tokens in a sequence. Based on a latency budget, retained layer ids are passed through by all tokens, ensuring KV Cache handling and retaining key layers. However, the limitation of this approach is that skipping is independent of the input sequence. In contrast, early exit strategies adapt layer skipping to the input sequence, offering more flexibility. In (Fan et al., 2019), a method akin to dropout randomly skips layers during training, but this leads to performance decline during the pre-fill stage. FFN-SkipLLM (Jaiswal et al., 2024) constrains skipping to FFN layers to avoid KV Cache issues but fails to fully exploit redundancy as discussed already. Song et al. (2024) is a very recent work that also explores greedily identifying layers to skip while preserving the model performance on a calibration dataset - however there are two major limitations of this work which are resolved in our paper - (A) first of all, the layer selection strategy is sequence independent although it can be made task-dependent by calibrating on task-specific datapoints - our approach for skipping layers is sequence dependent and is based on the input to a layer (B) SLEB does not explore the impact of fine-tuning on the layers to be skipped. On the other hand, our skipping strategy incorporates the trained router already - intuitively, the knowledge of skipping is transferred during finetuning. Chen et al. (2024) is another recent work that compresses models by identifying redundant layers - this is done by computing the average similarity between input/output pairs of a layer. However, as outlined in Song et al. (2024), such an approach suffers from the limitation that it does not take into account the joint association between the layers while skipping multiple layers. Moreover, like (Jaiswal et al., 2024), this work is neither sequence dependent nor takes the final model predictions into account while identifying the layers to skip.
Model Compression and Quantization Aware Training:
Orthogonal approaches to explore the latency/memory-performance trade-off in Large Language Models aim to build smaller models that approximate the performance of larger ones with reduced memory and latency costs. Key techniques include: 1) compressing model parameters into fewer bits (Frantar et al., 2022; Lin et al., 2024; Lee et al., 2024; Saha et al., 2023); 2) pruning the network by removing components like attention heads or neurons based on heuristics (Frantar & Alistarh, 2023; Ma et al., 2023b); and 3) distilling the large model into a smaller, faster counterpart (Agarwal et al., 2023; Gu et al., 2024). For further details, we refer to the survey by (Zhu et al., 2023). A significant body of work (Dettmers et al., 2024; Liu et al., 2023b; Peri et al., 2020; Li et al., 2023) has focused on quantization-aware training to reduce memory footprints and mitigate performance loss, starting with QLoRA (Dettmers et al., 2024). In a similar vein, our work proposes fine-tuning router-augmented models to improve layer skipping and reduce performance degradation, as pre-trained models do not account for layer skipping, leading to higher degradation with vanilla skipping.
Network Pruning:
Another orthogonal approach to improve the inference speed-up is to prune redundant network weights by zeroing them out. There has been a significant body of work on pruning model weights Frantar et al. (2022); Frantar & Alistarh (2023); Sun et al. (2022); Zhang et al. (2023) - most of these works can be categorized into two clusters namely unstructured pruning and structured pruning. In case of unstructured pruning, there is no structure to the inserted zeros and achieving speedups with modern GPU hardware tailored towards dense matrix multiplication is challenging. In fact, more than 90% sparsity is typically required to achieve any significant speedup Wang (2020); Shi et al. (2020). Therefore, structured pruning which is more amenable to GPU hardware has become prominent (2:4 pruning and sub-channel pruning). However, realizing desired speedups through these techniques have been difficult Song et al. (2024). Moreover, several approaches for dynamically deleting entire rows or columns of weight matrices have been proposed Ma et al. (2023a); Ashkboos et al. (2024); Liu et al. (2023c) to retain dense matrices but two limitations remain - (A) hardware support is extremely limited for realizing speedup gains (B) extensive finetuning is necessary to align the sparsification with linguistic abilities - this is because, such pruning techniques were not observed by the model during pre-training. Finally, note that several prior works Tang et al. (2024); Oren et al. (2024); Xiao et al. (2023) have imposed (query aware/ query agnostic) sparsity in the KV cache matrices to speed up self-attention mechanism via clever selection of the critical tokens necessary from the KV cache.
A.2 Details of Datasets
Machine Translation: For translation tasks, namely English-to-Chinese and English-to-German, we employ the WMT development sets from 2017 to 2020 for training/fine-tuning following the methodology outlined in previous studies (Liu et al., 2023a; Jiao et al., 2023). Translation performance is evaluated using the test set from the WMT 2022 dataset (Kocmi et al., 2022) which was developed using recent content from diverse domains. These domains include news, social media, e-commerce, and conversational contexts.
(Details in Appendix: A.5, Table: 8). Summarization: We use the popular CNN-DailyMail (CNN/DM) (Hermann et al., 2015) dataset which is a large collection (over 300k) of text summarization pairs, created from CNN and Daily Mail news articles. Each datapoint in this dataset comprises of an article (the body of the news article with 683 words on average) and the corresponding highlights (article summary as written by the article author). While the training set contains more than 287k samples, we have randomly chosen 4k samples for training both routers and LoRA. During training in our framework, the number of trainable parameters is small in both phases - therefore a small subset of data points is sufficient for training.
Inference is performed on the standard test set with 11,490 samples.
Question Answering: We use the popular Stanford Question Answering Dataset (SQuAD v1.0) (Rajpurkar et al., 2016), a widely-used benchmark for machine Question Answering. The dataset consists of over 100k question-answer pairs posed by crowd-workers on a set of over 500 Wikipedia articles. Each sample comprises a context (a passage from a Wikipedia article), a question (crafted to test comprehension of the passage) and the corresponding answer (a text span from the corresponding reading passage). Similarly to the CNN/DM dataset, 4k samples are chosen as random to train both routers and LoRA. The training and validation split contains 87,599 and 10,570 samples respectively. Evaluation is performed on the validation set Schuster et al. (2022) as the test set labels are not publicly released.
A.3 Evaluation Metrics
Quality-Based Metrics for Translation task:
  • BLEU Score: BLEU (Bilingual Evaluation Understudy) scores are used to measure the quality of translations. BLEU compares n-grams of the candidate translation to n-grams of the reference translation, providing a score between 0 and 1, with higher scores indicating better translations. In this evaluation, NLTK BLEU is employed, focusing on BLEU-1 and BLEU-2 scores.
  • COMET: COMET (Cross-lingual Optimized Metric for Evaluation of Translation) is used to assess translation quality further. COMET evaluates translations using a model trained to correlate well with human judgments. Specifically, Unbabel/XCOMET-XL  [^2] https://github.com/Unbabel/COMET is used in this evaluation. COMET provides a more nuanced assessment of translation quality by considering the intricacies of both source and target languages, beyond the n-gram matching used in BLEU.
Quality based Metrics for Summarization Task:
  • BERTScore: This metric quantifies semantic similarity between texts by leveraging contextual word embeddings.
    BERTScore captures meaning-based similarity rather than relying on exact word matches, providing a nuanced evaluation of text generation quality.
  • ROUGE: (Recall-Oriented Understudy for Gisting Evaluation) is a common metric - ROUGE-1 refers to overlap of unigrams between the system summary and reference summary. Similarly, ROUGE-L measures longest matching sequence of words.
Quality based Metrics for Question Answering Task:
  • Exact Match: This metric measures the percentage of predictions that exactly match the ground truth answer.
  • F1 score: Since EM is a highly stringent metric, we also report the F1 score which provides a more flexible evaluation of answer prediction. This metric also takes into account near-matches.
A.4 Training and Inference Setup
  • Training settings: We perform extensive experiments on two models, namely LLaMA-3-8B and LLaMA-3.2-3B from Meta, which consist of 32 and 28 layers, respectively. Training of routers and LoRA adapters is conducted on A100 80GB GPUs, with training/inference is performed in full precision to avoid performance degradation due to quantization. The training process employs our custom loss function and continues for a fixed number of epochs, terminating when the validation loss fails to improve over 4 consecutive steps. The learning rate is set between $1e^{-4}$ and $3e^{-4}$ a cosine scheduler is used to adjust the learning rate. Gradients are accumulated after 5 steps and the regularization coefficient $\lambda$ is fixed at 0.01. For LoRA fine-tuning, we employ a rank of 8, a dropout rate of 0.1, and a scaling factor (lora_alpha) of 32. We set the non-skipping penalization regularizer hyper-parameter $\beta=\alpha/3$ after tuning to maintain skipping level - where $\alpha$ is the non-skipping penalization hyperparameter used in the first phase and set according to the target speedup ratio. For translation, the maximum sequence length is set to 128 for router training and 256 for LoRA training. Similarly, for summarization, the maximum sequence length is set to 500 and 700 respectively. For Question Answering, this length is set to 512. Prompts for the different tasks regarding training/inference are shown in Appendix A.6.
  • Inference settings: For all the tasks, we set the temperature to 0.8 and enable top-k sampling over 10 tokens. The maximum number of tokens to be generated is set to 80 for WMT, 200 for CNN/DM and 32 for SQuAD. Caching is turned on during inference.
A.5 Training and testing split
WMT Summarization RC
Eng-to-German Eng-to-Chinese CNN/DM SQuAD
Train 3505 8983 3400 3400
Validation 876 998 600 600
Test 2038 2038 11490 10570
Table 8: Train-Validation-Test split for WMT, CNN/DM and SQuAD datasets
Figure 3: Comparison of LLaMA-3.2-3B (left) and LLaMA-3-8B (right) at 15% skipping rate on English-to-German Machine Translation Task. The graph shows how different layers contribute to the skipping behavior for the same dataset. Layers with no skipping, indicated by a 0% skipping rate, are not represented in the plot.
Figure 4: Comparison of LLaMA-3.2-3B (left) and LLaMA-3-8B (right) at 15% skipping rate on English-to-Chinese Machine Translation Task. The graph shows how different layers contribute to the skipping behavior for the same dataset. Layers with no skipping, indicated by a 0% skipping rate, are not represented in the plot.
Figure 5: Comparison of LLaMA-3.2-3B (left) and LLaMA-3-8B (right) at 15% skipping rate on CNN Summarization Task. Layers with no skipping, indicated by a 0% skipping rate, are not represented in the plot.
Figure 6: Comparison of LLaMA-3.2-3B (left) and LLaMA-3-8B (right) at 15% skipping rate on SQuAD Question-Answering Task. Layers with no skipping, indicated by a 0% skipping rate, are not represented in the plot.
A.6 Prompt Details
The prompt structures used for both training and inference are as follows:
  • For the machine translation task (English-to-German or English-to-Chinese), the following general prompt structure is used to train the routers and during final inference:
    ### Instruction:
        Translate the following sentences from English
    to German.
    ### Input:
        {Text to be translated}
    ### Response:
  • For the summarization task (used in CNN/DailyMail dataset), the prompt structure used is:
    ### Instruction:
        Summarize the news article in around 100-200
    words.
    ### Input:
        {Article to be summarized}
    ### Response:
  • For the Question Answering task (used in SQuAD dataset), the following prompt structure is utilized:
    ### Instruction:
        Answer the question based on the given passage.
    ### Passage:
        {context}
    ### Question:
        {Question to be answered}
    ### Response:
[…]
A.7 Layer-wise Skipping Statistics
Table 15 present the fraction of sequences that skip a particular block during the task for the LLaMA-3-8B model. If the corresponding cell in a row shows a value of 80.00, it implies that 80% of the sequences skip this block. It is important to note that the decision regarding which block to skip varies across different datasets and tasks. Additionally, partial skipping in some blocks, with varying percentages, suggests that while some sequences consider the layer important, others do not and therefore skip it during the decoding phase.
Figures 3, 4, 5, and 6 depict the distribution of skipped blocks when the model is configured to skip approximately 15% of the layers. These plots demonstrate the variation in block skipping based on the task and dataset, highlighting the differences in block importance. The bar plots further show that the determination of important layers depends not only on the dataset but also on the model architecture. For instance, Figure 3 illustrates that layers 7 and 9 are entirely skipped in LLaMA-3-8b, while they are rarely skipped in LLaMA-3.2-3b.
A.8 Detailed Result Table
Tables 9 and 11 present the detailed results for the LLaMA-3.2-3B model, while Tables 10 and 12 summarize the performance of the LLaMA-3-8B model for an additional skipping percentage on Machine Translation Task. The results are reported using BLEU (BLEU-1, BLEU-2) and COMET metrics, highlighting performance across different skipping percentages. Similarly, Table 13 presents cumulative results for both models reporting BERT F1, ROUGE-1 and ROUGE-L. Lastly, Table 14 presents Exact Match (EM) and F1 scores for both models for three skipping percentage variations.
Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 37.57 17.43 89.72
Base 31.19 13.67 81.66
15 Skip Decode Router + LoRA 23.24 11.07 44.58
Router 9.93 3.89 32.74
Random Skip Router + LoRA 18.99 4.78 47.26
Router 12.11 2.77 36.30
Unified Skip Router + LoRA 21.65 5.93 44.72
Router 16.41 3.99 39.81
FiRST (Ours) Router + LoRA 24.68 7.13 60.29
Router 16.47 4.11 43.04
25 Skip Decode Router + LoRA 16.85 7.62 32.33
Router 9.14 3.51 27.64
Random Skip Router + LoRA 8.95 1.07 30.97
Router 5.32 0.69 27.22
Unified Skip Router + LoRA 15.14 2.65 39.81
Router 10.21 1.62 30.86
FiRST (Ours) Router + LoRA 18.89 4.06 45.38
Router 9.51 1.44 29.08
35 Skip Decode Router + LoRA 1.62 0.36 23.30
Router 1.71 0.27 19.34
Random Skip Router + LoRA 4.72 0.15 25.08
Router 1.96 0.06 21.18
Unified Skip Router + LoRA 1.18 0.05 19.65
Router 0.90 0.03 20.36
FiRST (Ours) Router + LoRA 9.47 1.29 27.45
Router 5.74 0.56 25.03
Table 9: Machine Translation Results for English to German on LLaMA-3.2-3B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 41.78 21.74 93
Base 37.17 18.57 87.13
15 Skip Decode Router + LoRA 23.04 10.52 55.62
Router 3.99 1.18 23.33
Random Skip Router + LoRA 30.43 10.98 66.25
Router 26.54 8.77 60.27
Unified Skip Router + LoRA 28.92 10.64 59.34
Router 23.23 7.85 59.26
FiRST (Ours) Router + LoRA 38.01 17.89 82.14
Router 28.83 11.8 67.74
25 Skip Decode Router + LoRA 13.67 6.00 31.47
Router 3.24 0.92 21.55
Random Skip Router + LoRA 6.01 0.91 29.71
Router 3.65 0.49 29.95
Unified Skip Router + LoRA 15.67 3.36 31.69
Router 12.58 2.65 32.15
FiRST (Ours) Router + LoRA 17.84 4.14 34.95
Router 9.67 1.37 26.01
35 Skip Decode Router + LoRA 5.55 0.33 23.85
Router 3.03 0.82 20.03
Random Skip Router + LoRA 1.80 0.12 25.56
Router 1.44 0.06 25.34
Unified Skip Router + LoRA 6.44 0.77 22.05
Router 3.92 0.51 22.88
FiRST (Ours) Router + LoRA 6.39 0.42 19.96
Router 3.7 0.14 21.41
Table 10: Machine Translation Results for English to German on LLaMA-3-8B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 51.81 30.04 79.13
Base 32.10 17.92 61.84
15 Skip Decode Router + LoRA 38.14 21.68 46.70
Router 9.59 4.84 34.14
Random Skip Router + LoRA 38.75 17.39 57.79
Router 13.56 5.64 35.86
Unified Skip Router + LoRA 36.96 17.35 57.10
Router 22.43 9.37 45.16
FiRST (Ours) Router + LoRA 45.66 23.66 67.45
Router 22.69 10.92 54.55
25 Skip Decode Router + LoRA 27.81 15.74 42.01
Router 7.01 3.41 29.51
Random Skip Router + LoRA 25.12 9.85 45.53
Router 10.07 3.69 31.50
Unified Skip Router + LoRA 30.56 12.27 42.30
Router 15.21 5.51 31.63
FiRST (Ours) Router + LoRA 32.92 13.74 41.66
Router 10.46 3.55 27.83
35 Skip Decode Router + LoRA 2.84 1.10 22.30
Router 2.36 1.04 18.32
Random Skip Router + LoRA 4.17 1.52 32.81
Router 2.11 0.76 23.63
Unified Skip Router + LoRA 3.26 0.52 19.12
Router 2.20 0.37 19.82
FiRST (Ours) Router + LoRA 20.01 6.96 28.10
Router 6.54 1.78 23.13
Table 11: Machine Translation Results for English to Chinese on LLaMA-3.2-3B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
Skipping (%) Model Type Configuration BLEU-1 BLEU-2 COMET
0 Original Model Base + LoRA 56.94 35.56 82.66
Base 38.02 22.46 68.95
15 Skip Decode Router + LoRA 28.73 15.84 55.98
Router 4.74 2.33 21.75
Random Skip Router + LoRA 47.88 25.75 67.32
Router 36.60 18.66 59.89
Unified Skip Router + LoRA 46.61 25.01 69.58
Router 27.28 13.35 54.57
FiRST (Ours) Router + LoRA 48.35 26.57 68.63
Router 17.55 8.68 42.76
25 Skip Decode Router + LoRA 20.03 10.85 33.85
Router 3.78 1.84 20.93
Random Skip Router + LoRA 11.69 4.66 27.73
Router 7.37 2.81 35.16
Unified Skip Router + LoRA 34.90 15.75 50.59
Router 17.74 7.35 38.74
FiRST (Ours) Router + LoRA 35.79 15.66 56.92
Router 11.01 3.23 25.45
35 Skip Decode Router + LoRA 12.24 6.27 25.23
Router 3.64 1.74 22.84
Random Skip Router + LoRA 7.38 2.19 27.35
Router 4.47 1.12 29.46
Unified Skip Router + LoRA 7.51 2.10 20.25
Router 3.87 1.06 21.24
FiRST (Ours) Router + LoRA 15.66 3.95 26.80
Router 6.13 1.54 22.89
Table 12: Machine Translation Results for English to Chinese on LLaMA-3-8B: BLEU-1, BLEU-2 and COMET scores for various skipping strategies.
Skip (%) Model Type BERT R-1 R-L
0 Original Model wLoRA 84.87 28.46 16.99
Base 82.29 23.49 14.66
15 Skip Decode R + LoRA 84.74 22.04 17.54
Router 82.53 13.68 9.30
Random Skip R + LoRA 83.70 24.60 15.01
Router 81.10 19.64 13.07
Unified Skip R + LoRA 84.25 24.35 14.3
Router 80.3 16.61 10.95
FiRST (Ours) R + LoRA 85.14 31.8 20.13
Router 81.25 20.2 13.01
20 Skip Decode R + LoRA 82.57 20.41 14.87
Router 81.62 13.48 9.19
Random Skip R + LoRA 81.39 21.57 13.83
Router 79.23 15.51 10.93
Unified Skip R + LoRA 82.93 22.3 13.37
Router 80.32 16.51 11.15
FiRST (Ours) R + LoRA 82.8 27.65 17.84
Router 79.32 16.28 10.85
27 Skip Decode R + LoRA 79.92 10.67 10.32
Router 77.27 9.59 7.00
Random Skip R + LoRA 76.40 11.45 7.89
Router 77.45 12.56 9.08
Unified Skip R + LoRA 80.28 15.94 9.89
Router 77.43 10.97 7.68
FiRST (Ours) R + LoRA 77.5 14.65 10.45
Router 75.6 9.39 6.92
Skip (%) Model Type BERT R-1 R-L
0 Original Model wLoRA 84.89 28.37 17.02
Base 71.85 19.34 12.00
15 Skip Decode R + LoRA 83.20 21.71 13.74
Router 80.97 9.74 6.87
Random Skip R + LoRA 79.52 20.18 12.10
Router 68.20 10.10 7.10
Unified Skip R + LoRA 81.53 18.89 11.72
Router 70.01 12.49 8.68
FiRST (Ours) R + LoRA 83.17 26.47 16.79
Router 70.98 16.47 10.51
24 Skip Decode R + LoRA 78.55 15.83 6.74
Router 76.91 13.29 8.86
Random Skip R + LoRA 80.00 16.33 10.07
Router 67.88 8.49 5.96
Unified Skip R + LoRA 79.31 15.88 10.69
Router 68.86 9.17 6.97
FiRST (Ours) R + LoRA 80.25 21.28 13.89
Router 69.17 12.36 8.43
28 Skip Decode R + LoRA 70.69 8.76 6.74
Router 40.99 2.05 1.23
Random Skip R + LoRA 79.45 14.69 9.23
Router 67.48 8.14 5.64
Unified Skip R + LoRA 78.57 11.74 7.47
Router 68.23 8.12 5.66
FiRST (Ours) R + LoRA 77.48 15.98 11.14
Router 67.14 8.09 6.00
Table 13: Quality Analysis on Summarization (CNN/DM dataset) on LLaMA-3-8B (left) and LLaMA-3.2-3B (right): BERT F1, Rouge-1 and Rouge-L scores are reported for varying skipping levels. Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning. FiRST with fine-tuning, improves upon Unified Skipping for all skipping levels on both Rouge-1 and Rouge-L and is competitive on BERT F1.
Skip (%) Model Type EM F1
0 Original Model wLoRA 73.93 85.99
Base 19.46 36.73
10 Skip Decode R + LoRA 60.14 65.33
Router 16.38 31.48
Random Skip R + LoRA 65.73 80.08
Router 18.25 33.75
Unified Skip R + LoRA 55.54 74.58
Router 17.39 32.91
FiRST (Ours) R + LoRA 70.85 83.61
Router 14.58 31.52
20 Skip Decode R + LoRA 45.00 55.10
Router 10.68 26.69
Random Skip R + LoRA 47.79 66.37
Router 6.71 22.46
Unified Skip R + LoRA 52.87 69.28
Router 18.18 32.51
FiRST (Ours) R + LoRA 60.60 75.49
Router 13.21 27.48
30 Skip Decode R + LoRA 30.77 48.38
Router 10.67 28.52
Random Skip R + LoRA 25.45 42.68
Router 3.55 15.49
Unified Skip R + LoRA 25.61 38.55
Router 15.19 28.11
FiRST (Ours) R + LoRA 38.20 52.68
Router 3.64 13.30
Skip (%) Model Type EM F1
0 Original Model wLoRA 73.07 84.17
Base 18.92 37.74
10 Skip Decode R + LoRA 60.79 75.00
Router 20.00 31.55
Random Skip R + LoRA 64.78 77.27
Router 13.76 28.59
Unified Skip R + LoRA 65.03 77.53
Router 13.16 32.31
FiRST R + LoRA 69.44 81.35
Router 12.79 28.37
20 Skip Decode R + LoRA 40.12 40.00
Router 20.45 37.62
Random Skip R + LoRA 11.32 38.34
Router 6.75 15.51
Unified Skip R + LoRA 37.39 52.49
Router 7.81 18.20
FiRST R + LoRA 39.70 54.59
Router 5.52 15.33
30 Skip Decode R + LoRA 20.68 23.08
Router 0.78 3.83
Random Skip R + LoRA 0.30 8.87
Router 0.62 5.88
Unified Skip R + LoRA 7.39 13.72
Router 0.37 6.15
FiRST R + LoRA 33.99 50.37
Router 2.55 10.26
Table 14: SQuAD performance on LLaMA-3-8B (left) and LLaMA-3.2-3B (right): EM (Exact Match) and F1 scores are reported for varying skipping levels. Note that R + LoRA corresponds to Router Augmentation followed by LoRA fine-tuning (in the proposed FiRST framework) and wLoRA stands for Base Model with LoRA fine-tuning.
Layer $\downarrow$ R R+L R R+L R R+L
$\alpha\rightarrow$ 0.005 0.005 0.01 0.01 0.025 0.025
0 0.00 0.00 0.00 0.00 0.00 0.00
1 0.00 0.00 0.00 0.00 0.00 0.00
2 0.00 0.00 0.00 0.00 0.00 0.00
3 0.00 0.00 0.00 0.00 0.00 0.00
4 0.00 0.00 0.00 0.00 0.00 0.00
5 0.00 0.00 0.00 0.00 0.00 0.00
6 0.00 0.00 0.00 0.00 0.00 0.00
7 100.00 100.00 100.00 100.00 100.00 100.00
8 100.00 100.00 100.00 100.00 100.00 100.00
9 100.00 100.00 0.00 0.00 0.10 1.32
10 0.00 0.00 0.00 0.00 89.25 72.67
11 0.00 0.00 0.00 0.00 0.00 0.00
12 0.00 0.00 100.00 100.00 100.00 100.00
13 0.00 0.00 0.00 0.00 0.00 0.00
14 0.00 0.00 0.00 0.59 0.00 0.25
15 0.00 0.00 100.00 99.95 100.00 99.36
16 0.00 0.00 0.10 2.45 100.00 99.95
17 0.00 0.00 0.00 0.00 0.00 0.00
18 97.79 32.24 100.00 100.00 100.00 100.00
19 0.00 0.00 99.85 91.17 100.00 99.61
20 0.00 0.00 100.00 100.00 100.00 99.46
21 99.85 98.72 100.00 100.00 100.00 100.00
22 0.00 0.00 0.00 0.00 0.00 0.00
23 0.00 0.00 24.14 0.79 99.75 91.66
24 0.00 0.00 0.00 0.00 0.00 0.00
25 0.00 0.00 0.00 0.00 0.00 0.00
26 0.00 0.00 57.31 2.45 100.00 86.02
27 0.00 0.00 0.00 0.00 0.00 0.00
28 0.00 0.00 0.00 0.00 0.00 0.05
29 0.00 0.00 0.00 0.00 0.00 0.00
30 0.00 0.00 0.00 0.00 0.00 0.00
31 0.00 0.00 0.00 0.00 0.00 0.00
Avg 15.55 13.47 27.54 24.92 37.16 35.95
Table 15: Variation in skipping percentage (15-35%) with the Non-skip Penalization Loss coefficient $\alpha$ for LLaMA-3-8B on Machine Translation (English-to-German). As $\alpha$ increases, the skipping percentage also increases for both the Router-Only and Router-Augmented LoRA fine-tuned models. Similar trends are observed for other datasets as well.